OpenStudy (snowystorm777):

Find the value of the expression 4a^4 − 2b^2 + 40 when a = 2 and b = 7

pooja195 (pooja195):

Hi welcome to openstudy ^-^ So basically plug in the numbers given \[\huge~\rm~\bf 4(2)^4-2(7)^2\]

OpenStudy (snowystorm777):

Thanks,so it's 16-98?

pooja195 (pooja195):

hmm nope \[\huge~\rm~\bf 4^2=? \]

OpenStudy (snowystorm777):

Wait how did u get that?

pooja195 (pooja195):

When we plug in the given values we get: \[\huge~\rm~\bf 4(2)^4-2(7)^2\] First lets get whole numbers 2 x 2 x 2 x 2 =?

OpenStudy (snowystorm777):

16

pooja195 (pooja195):

Good :) 7^2=7 x 7 =?

OpenStudy (snowystorm777):

14

pooja195 (pooja195):

No we are multiplying 7 times 7=?

OpenStudy (snowystorm777):

oops 49

pooja195 (pooja195):

Good :) 16 x 4=64 \[\huge~\rm~64−(2)(49)\]

OpenStudy (snowystorm777):

I keep getting 3038 I dont know why??

pooja195 (pooja195):

49 x 2 =?

rvc (rvc):

where r u stuck dear?

rvc (rvc):

u got 49 correct? @snowystorm777

pooja195 (pooja195):

I'll let @rvc take over :) I dont want to confuse you

rvc (rvc):

i would like u to wait @pooja195 till we complete :)

OpenStudy (snowystorm777):

ok

rvc (rvc):



OpenStudy (snowystorm777):

I keep getting 3038???

OpenStudy (snowystorm777):

U there??

rvc (rvc):

lets do 49 X 2 first

OpenStudy (snowystorm777):

98

rvc (rvc):

cool now 64-98=?

OpenStudy (snowystorm777):

-34

rvc (rvc):

34+40 =?

OpenStudy (snowystorm777):

74

rvc (rvc):

see i made a mistake

rvc (rvc):

it should be -34 isnt it?

OpenStudy (snowystorm777):

yah I wrote -34

rvc (rvc):

-34+40 =?

OpenStudy (snowystorm777):

6

rvc (rvc):

bravo thanks to pooja195 u r correct

OpenStudy (snowystorm777):

Thanks!